OpenStudy (anonymous):

2sin^2x - 5sinx - 3=0 Part 1: Rewrite the equation by substituting the expression u in for sin x. Part 2: Factor the quadratic expression. Rewrite the equation with factors instead of the original polynomial. Part 3: Use the zero product property to solve the quadratic equation. Part 4: Rewrite your solutions to part 3 by replacing u with sin x. Part 5: Solve the remaining equations for x, giving all solutions to the equation.

OpenStudy (campbell_st):

well this is an equation that can be rewritten as a quadratic. making the required substitution gives \[2u^2 -5u - 3 = 0\] so now you are solving the quadratic

OpenStudy (campbell_st):

so next you need to factor the quadratic...

OpenStudy (anonymous):

how would i solve that?

OpenStudy (campbell_st):

well here is my method for a quadratic \[ax^2 + bx + c = 0\] multiply a and c so in your question you need to multiply 2 and -3 then find the factors of the answer that add to -5... the larger factor is negative. Any thoughts..?

OpenStudy (anonymous):

So for the factors do I just pick any numbers that add up to -5, like -10+5?

OpenStudy (campbell_st):

nope if you multiply 2 and -3 you get -6 you need to find the factors of -6 that add to -5... the larger is negative

OpenStudy (anonymous):

oh so like -6+1

OpenStudy (campbell_st):

thats it,,, next part to factoring write it as (ax + factor 1)(ax + factor 1) ------------------------- = 0 a so you get (2u -6)(2u + 1) --------------- = 0 2

OpenStudy (campbell_st):

then remove the common factor in the 1st binomial 2(u - 3)(2u + 1) ---------------- = 0 2 so the factored form is (u -3)(2u + 1) = 0 now you need to solve it

OpenStudy (anonymous):

would it be 2u^2-5u-3=0 ?

OpenStudy (campbell_st):

now all that has been done is to write the original equation in factored form. As required by Part 2. Part 3 let u - 3 = 0 2u + 1 = 0 solve both equations for u....

OpenStudy (anonymous):

so u=3 and u= -1/2

OpenStudy (anonymous):

I still need help with part 5